THE NEAR-IR HIGH RESOLUTION SPECTRUM OF JET-COOLED ETHYL PEROXY RADICALS

\underline{PATRICK DUPRÉ}, PATRICK RUPPER and TERRY A. MILLER, Department of Chemistry, The Ohio State University, 120 W 18^{th} Ave, Columbus OH, 43210.

The jet-cooled high resolution CRDS spectra of the ethyl peroxy radicals, C_{2}H_{5}OO and C_{2}D_{5}OO have been recently obtained for the Ã^{2}A′ ← X̃^{2}A′′ 0_{0}^{0} band. The observation of this weak transition: signal of a few ppm/pass for the most populated G conformer has been possible by using a narrow linewidth (Fourier-transform-limited) laser source^{a} and a hydrogen Raman shifter (resolution: 250 − 300 MHz). The radicals are produced by creating a plasma (by using a DC or a RF discharge) inside a pulsed slit nozzle^{b}. We will discuss the analysis of these complex spectra.

aP. Dupré and T.A. Miller, Rev. Sci. Instrum. 78, 033102 (2007)

bS. Wu, P. Dupré and T.A. Miller, Phys. Chem. Chem. Phys. 8, 1682 (2006)